OpenStudy (anonymous):

Write an equation and a solution for the problem. A health food store sells oatmeal for $3.60 per pound and bran flakes for $4.80 per pound. How many pounds of each should be used to get a mixture of 30 pounds that sells for $4.00 a pound?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

20 pounds of oatmeal are needed 10 pounds of bran flakes are needed

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

Write an equation & a solution for it.

OpenStudy (anonymous):

Let a = pounds of oatmeal needed Let b = pounds of bran flakes needed given: (1) a + b = 30}}} {{{ (360a + 480b) / 30 = 400 }}} (in cents) {{{ 360a + 480b = 12000 }}} (2) {{{ 3a + 4b = 100 }}} Multiply both sides of (1) by {{{3}}} and subtract (1) from (2) (2) {{{ 3a + 4b = 100 }}} (1) {{{-3a - 3b = -90}}} {{{b = 10}}} and, since (1) {{{a + b = 30}}} (1) {{{a = 30 - 10}}} (1) {{{a = 20}}} 20 pounds of oatmeal are needed 10 pounds of bran flakes are needed check: {{{ (360a + 480b) / 30 = 400 }}} {{{ (360*20 + 480*10) / 30 = 400 }}} {{{ (7200 + 4800) / 30 = 400 }}} {{{ 12000/30 = 400 }}} {{{ 12000 = 12000 }}}

OpenStudy (anonymous):

got ?

OpenStudy (anonymous):

Thank you. (:

OpenStudy (anonymous):

Will you help me with my other question?